# SCAR: Scheduling Multi-Model AI Workloads on Heterogeneous Multi-Chiplet Module Accelerators

### v1

@@ Abstract
Therefore, in this work, we explore the opportunity in the heterogeneous dataflow MCM AI accelerators. We identify the scheduling of multi-model workload on heterogeneous dataflow MCM AI accelerator is an important and challenging problem due to its significance and scale, which reaches O($10^{18}$) scale even for a single model case on 6x6 chiplets. We develop a set of heuristics to navigate the huge scheduling space and codify them into a scheduler with advanced techniques such as inter-chiplet pipelining. Our evaluation on ten multi-model workload scenarios for datacenter multitenancy and AR/VR use-cases has shown the efficacy of our approach, achieving on average 35.3% and 31.4% less energy-delay product (EDP) for the respective applications settings compared to homogeneous baselines.
@@ I Introduction
Researchers have actively explored the MCM for AI, focusing on the dataflow mapping (i.e., loop ordering, parallelization, and tiling) and workload orchestration onto chiplets considering the network-on-package (NoP) and other communication constraints [48, 52, 55, 40]. For example, Simba [48] proposed a scalable MCM inference architecture that enables chiplets to either act as standalone inference engines or collaborate as groups for a layer. Although such works have successfully delivered promising performance and energy efficiency than monolithic designs, they mostly focused on single-model workloads targeting homogeneous chiplets. Unlike single-model workloads, multi-model workloads introduce major challenges to such homogeneous MCMs because of the ML operator heterogeneity (e.g., operator types and tensor sizes) and resulting diverse dataflow preferences [26]. Also, multi-model workloads often involve model level dependency and concurrency [27], which adds complex considerations to the scheduling problem.
@@
Therefore, considering the new trend with multi-model AI workloads such as multi-tenancy [13, 56, 28] and AR/VR [27], we propose to explore heterogeneous chiplet-based MCM with AI accelerator chiplets with various dataflows to address the workload heterogeneity and concurrency. We consider inter-layer pipelining to enhance in-package data reuse and reduce offchip traffic. We formulate the scheduling problem and develop effective heuristics to navigate the huge scheduling space, whose problem scale is as big as O($10^{18}$) on a 6x6 chiplet MCM AI accelerator system even running a single model (BERT-L).
We evaluate five MCMs including three heterogeneous MCMs on ten multi-model scenarios: the first five scenarios are curated using MLPerf [46] to represent datacenter multi-tenancy scenarios. The models are selected based on recent datacenter model usage trends [13, 19] and the trend of large language model adoptions (e.g., GPT-L [44]), future-proofing emerging AI workloads such as AI assistant [35]. The other five scenarios are curated for AR/VR usage scenarios from XRBench as a practical use case for edge multi-model workloads  [27]. The evaluation results show the effectiveness of heterogeneous MCM combined with our scheduling method. Compared to the homogeneous MCM [48] running NVDLA [39] and Shi-diannao [8] style dataflows, heterogeneous MCM, on average, achieved 35.3% and 31.4% less energy-delay product (EDP) in each domain, respectively. Moreover, we performed ablation studies on the efficacy of our greedy scheduling algorithm, and found that it leads to superior multi-model schedules compared to other approaches, achieving 21.8% and 8.6% execution speedups and energy improvement on heterogeneous MCMs. We summarize our contributions as follows:
@@
    We propose to explore heterogeneous dataflow MCM for emerging AI workloads with multiple models running concurrently for the first time.
@@
    Based on the formulation, we develop a scheduler that thoroughly considers heterogeneous MCM and multi-model workloads. The scheduler employs advanced scheduling techniques, such as inter-layer pipelining, dynamic chiplet regrouping utilizing latest representation such as the resource allocation tree [4].
@@
    We analyze the costs and benefits of heterogeneous dataflow MCM using future-proof multi-model workloads motivated by recent industry use cases and present the importance of the scheduling problem.
II Background
We discuss examples of emerging multi-model AI workloads and chiplet-based MCM AI accelerators.
@@ II-A Multi-model AI Workloads
The success of AI algorithms in individual tasks (e.g., hand tracking, depth estimation, and speech recognition) led to the emergence of multi-model AI workloads, which include multi-tenant workloads at data centers [13, 56, 28] and real-time multi-model workloads such as AR/VR [27]. We summarize example multi-model AI workloads from industrial use cases in Table II. The models in such workloads are diverse in terms of the tasks and input modalities. For example, an industrial data center multi-tenant AI workload suite [13] includes a face recognition model based on support vector machine, recommendation models based on multi-layer perceptron, and a speech recognition model based on recurrent neural network (RNN). More recent workloads in data center AI workload include large language models [36], which adds more heterogeneity to the multi-model AI workloads. As discussed in prior works [26, 27], such multi-model workloads involve high heterogeneity in AI operators (or layers), which is one of the major challenges to accelerators that specialize the architecture and dataflow for a specific set of workloads.
II-B MCM AI Accelerators
Multi-chip Modules (MCM) comprise small functional dies (chiplets) that are packaged together to build a larger system. Chiplets are interconnected via on-package links typically through silicon interposer or organic substrates to create a network-on-package (NoP) [54, 20, 2]. A typical chiplet, in the context of a DNN accelerator, comprises off-chip memory, a global shared memory, and an array of processing elements connected via a Network-on-Chip (NoC) [6]. Advantages of the chiplet-based MCM architecture include the modularity and scalability to systems of varied scales simply by adjusting the number of chiplets placed on the package as well as low verification cost [38]. Based on such benefits, many chiplet-based MCMs have been developed for scalable DL inference [48, 5, 40, 51]. Such MCM accelerators successfully scaled up the systems up to 256 chiplets with 1 million processing engines  [40]. However, the effectiveness of a chiplet-based MCM system heavily depends on the careful distribution of computation amongst the different chiplets while balancing the added NoP/NoC communication costs.
II-C Scheduling space
As discussed in previous works [41, 26, 4], scheduling AI workloads on an accelerator can be considered as assigning a set of computations in various granularity (e.g., layer or compute tile) to each compute unit and ordering the computation. That is, the scheduling process is spatially and temporarily partitioning a workload onto a target accelerator architecture [4]. However, with that formulation, the scheduling space of multi-model workloads onto shared MCM accelerators is intractably large and high-dimensional, as discussed in Section I.
@@ II-C Scheduling space
One approach to address the complexity is formulating the problem into multi-level decision problem where each decision subspace is a tractable problem [26, 6].We adopt a similar approach and formulate the multi-model workload scheduling on MCM as multiple-level decision problem, as shown in Figure 2. We discuss details of our problem formulation in detail and performance modeling methodology next.
@@ III System Modeling and Problem Formulation
To develop a systematic approach to navigate complex scheduling space, we formulate the scheduling problem of the multi-model workloads running on a heterogeneous MCM AI accelerator.
@@ III-A Base Formulation
We assume the 2D mesh topology for NoP like Simba [48], and chiplets on two sides (left and right) of the packages have off-chip interfaces.
@@ III-B Workload Partitioning Space
To develop a systematic optimization algorithm for layer segmentation within each time window, we need to define the conditions of valid layer segments. We define the condition as follows:
@@
$\bigcup_{sg\in SG}sg=L(tw(Sc))\land\forall sg_{i}\neq sg_{j}\in SG,sg_{i}\cap sg_{j}=\emptyset$
@@
$\bigcup_{tw\in TW(Sc)}tw=L(Sc)\land\forall tw_{i}\neq tw_{j}\in TW(Sc),tw_{i}\cap tw_{j}=\emptyset$
Both Theorem 1 and  Theorem 2 indicate that the time windows and segments need to be partitions of the workload and time window layer, respectively. Combining all definitions in this section, we formulate the workload partitioning space into the time window and segment as follows:
@@ III-C Scheduling Space
=$\text{=}\{(sg,c,j)|sg\in SG\land c\in C_{H}\land j\in\mathbb{N}\land valid(sg,tw(Sc))\}$
@@ III-D Scheduling Problem
The optimization metric can be chosen by users depending on the use case. In our scheduler, we adopt a comprehensive and customizable score that thoroughly consider all of latency, energy, and energy-delay product (EDP), allowing users to configure their own optimization metrics, which can the mentioned frequently used metrics or a user-defined function that takes a schedule instance and generates a custom metric.
III-E Latency Modeling
To develop a scheduler based on the scheduling problem formulation in Section III-C, we need to be able to evaluate each schedule on target MCM AI accelerator hardware. For that, we extend MAESTRO [24] to the chiplet domain and model the latency of MCM AI accelerators concurrently executing multi-model workloads on a shared MCM system in a bottom-up fashion. We discuss our latency evaluation methodology in detail, focusing on our extension for the MCM and multi-model workloads.
@@ III-E Latency Modeling
$Lat^{comp}(l)$ being the layer computation cost dependent on the AI accelerator chiplet parameters Definition 2; $Lat^{ip\_com}(l)$ is latency incurred from loading the layer operands (input activations and weights), and $Lat^{op\_com}(l)$ from transmitting the output activation to a subsequent layer. As for $Lat^{com}$, it is defined as:
@@
where assuming sufficient memory for double-buffering on each chiplet accelerator, communication costs become incurred when transmitting data to/from another chiplet on package or the offchip memory. The first term $\frac{Sz_{data}}{BW}$ reflects transmission latency; the second term is captures propagation latency across $n_{hops}$ between the source and destination; $\delta$ is an additional latency term for potential NoP traffic conflicts; $Lat_{mem}$ is the cost from read/write access of data at the offchip memory.
@@
The first term represents the sum of individual layer computational latencies; $Lat^{ip\_com}$ is the initial external data transfer loading costs of necessary off-chiplet input activation and parameter weights; $Lat^{op\_com}$ is the transmission latency from transmitting segment output data to the next segment or writing back to memory.
Based on the segment’s latency, the time window latency can be defined as:
$Lat(tw)=\max\limits_{SG_{m}\subset SG}\begin{cases}\sum_{sg_{k}\in SG_{m}}Lat(sg_{k}),\hfill\text{if end-to-end}\\ \max\limits_{sg_{k}\in SG_{m}}Lat(sg_{k}),\hfill\text{if pipelining}\end{cases}$
where $SG_{m}$ represents the set of segments in a time window associated with a model $m$. The latency of $SG_{m}$ can either be the maximum out of all $m$ segments if pipelining or their summation in the base case. The latency of the entire window is the max out of all $SG_{m}\subset SG$.
@@
III-F Energy Modeling
Albeit similar to latency, Energy costs are always aggregated. The base communication energy cost is defined as:
$E^{com}=\begin{cases}0,\hfill\text{if same chiplet}\\ Sz_{data}\times E_{tx\_bit}\times n_{hops},\hfill\;\;\text{if same package}\\ Sz_{data}\times E_{tx\_bit}\times n_{hops}+E_{mem},\;\hfill\text{if offchip}\end{cases}$
where the energy incurred from moving data across the package is equal to the product of data size, number of hops, and the per-bit transmission energy ($E_{tx\_bit}$). In case of an offchip data transmission, the cost of memory access $E_{mem}$ is added.
The overall energy consumption across the entire time windows can be computed as the sum of energies of its constituent components as follows:
$E(l)$ $=E^{ip\_com}(l)+E^{comp}(l)+E^{op\_com}(l)$
$E(tw)$ $=\sum_{sg_{k}\in SG}E(sg_{k}),\;\;\;E(Sc)=\sum_{tw_{j}\in TW}E(tw_{j})$
IV Scheduling Framework
Figure 2: An overview of our scheduling algorithm.

### v2

@@ Abstract
Therefore, in this work, we explore the opportunity in the heterogeneous dataflow MCM AI accelerators. We identify the scheduling of multi-model workload on heterogeneous dataflow MCM AI accelerator is an important and challenging problem due to its significance and scale, which reaches O($10^{56}$) even for a two-model workload on 6x6 chiplets. We develop a set of heuristics to navigate the huge scheduling space and codify them into a scheduler, SCAR, with advanced techniques such as inter-chiplet pipelining. Our evaluation on ten multi-model workload scenarios for datacenter multitenancy and AR/VR use-cases has shown the efficacy of our approach, achieving on average 27.6% and 29.6% less energy-delay product (EDP) for the respective applications settings compared to homogeneous baselines.
Index Terms:
AI accelerators, Multichip modules, Chiplets, Scheduling algorithms, Performance analysis.
@@ I Introduction
Researchers have actively explored the MCM for AI, focusing on the dataflow mapping (i.e., loop ordering, parallelization, and tiling) of each layer and workload orchestration onto chiplets considering the network-on-package (NoP) and other communication constraints [64, 68, 71, 55]. For example, Simba [64] proposed a scalable MCM inference architecture that enables chiplets to either act as standalone inference engines or collaborate as groups for a layer. Although such works have successfully delivered promising performance and energy efficiency than monolithic designs, they mostly focused on single-model workloads targeting homogeneous chiplets. Unlike single-model workloads, multi-model workloads introduce major challenges to such homogeneous MCMs because of the ML operator heterogeneity (e.g., operator types and tensor sizes) and resulting diverse dataflow preferences [37]. Also, multi-model workloads often involve model level dependency and concurrency [37, 38, 51, 34, 56], which adds complex considerations to the scheduling problem.
@@
Therefore, considering the new trend with multi-model AI workloads in industry, such as multi-tenancy [23, 72, 40] and AR/VR [38], we explore heterogeneous chiplet-based MCM with AI accelerator chiplets with various dataflows, as a future-proof option. To exploit the benefits of heterogeneous MCM accelerators, we consider inter-layer pipelining to enhance in-package data reuse and reduce offchip traffic. We formulate the scheduling problem and develop effective heuristics to navigate the huge scheduling space, whose problem scale is as big as O($10^{56}$) even for a two-model workload (ResNet-50 [24] and UNet [63]) on a 6x6 chiplet MCM AI accelerator system (as in Simba [64]).
We evaluate ten MCMs including seven heterogeneous MCMs on ten multi-model scenarios: the first five scenarios are curated using MLPerf inference benchmark [62] representing datacenter multi-tenancy scenarios. The models are selected based on recent datacenter model usage trends [23, 29] and the trend of language model adoptions (e.g., GPT-L [60]), future-proofing emerging AI workloads such as AI assistant [47]. The other five scenarios are curated for AR/VR usage scenarios from XRBench as a practical use case for edge multi-model workloads  [38].
The evaluation results show that heterogeneous MCM combined with our scheduling method is promising for heavy multi-model workloads, which is projected by recent trend. Compared to the homogeneous MCM [64] running NVDLA [52] and Shi-diannao [16] style dataflows, heterogeneous MCM, on average, achieved 27.6% and 29.6% less energy-delay product (EDP) in each domain, respectively. We also showcase that our scheduler can identify schedules that can reduce EDP to 0.3$\times$ that of single-model schedulers like NN-baton [68]. We summarize our contributions as follows:
@@
    As a promising future-proof architecture for heavy multi-model workloads, we explore heterogeneous dataflow MCM for emerging AI workloads with multiple models running concurrently.
@@
    Based on the formulation, we develop a scheduler that thoroughly considers heterogeneous MCM and multi-model workloads. The scheduler employs advanced scheduling techniques (inter-layer pipelining, dynamic chiplet regrouping) with resource allocation tree representation [8].
@@
    We analyze the costs and benefits of heterogeneous dataflow MCM using industry use case-inspired multi-model workloads and present the importance of the scheduling problem.
II Background and Motivation
@@ II-A Multi-model AI Workloads
Multi-model AI Workloads. The success of AI algorithms in individual tasks (e.g., hand tracking, depth estimation, speech recognition) led to the emergence of multi-model AI workloads, which include multi-tenant workloads at data centers [23, 72, 40] and real-time multi-model workloads such as AR/VR [38]. We summarize example multi-model AI workloads from industrial use cases in Table III. The models in such workloads are diverse in terms of the tasks and input modalities. For example, an industrial data center multi-tenant AI workload suite [23] includes a face recognition model based on support vector machine, recommendation models based on multi-layer perceptron, and a speech recognition model based on recurrent neural network (RNN). More recent workloads in data center AI workload include large language models [48], which adds more heterogeneity to the multi-model AI workloads. As discussed in prior works [37, 38], such multi-model workloads involve high heterogeneity in AI operators (or layers), which is one of the major challenges to accelerators that specialize the architecture and dataflow for a specific set of workloads.
II-B Scheduling AI workloads on AI Hardware and MCMs
Scheduling AI workloads considers the assignment of computations (e.g., model, layer, or tile) to target hardware platforms and their constituent computing units. We provide a brief on the state of AI workloads scheduling practices.
Scheduling on CPU/GPU systems. modern systems (servers) typically employ GPUs and/or CPUs for inference services [23, 12, 13, 22, 54, 65, 72, 76, 4]. Most of these computing units are based on homogeneous cores – or simple heterogeneity such as big and little cores in CPUs, or CUDA and Tensor cores in GPUs. Traditionally, the scheduling in such settings is concerned with the coarse assignment of models to computing units, leaving the operator assignments to be performed in a direct manner (e.g., all GEMM operations to Tensor Cores in a GPU). As multi-model workloads proliferated, new features (e.g., GPU sharing) emerged to improve inference services for small-batch inference tasks [12, 11]. Still, the limited programmer/compiler control and the cache-based memory systems restrict CPUs/GPUs from engaging multi-model workload scheduling on a finer granularity.
Scheduling on customized AI accelerators. Customized AI accelerators (Google’s TPU [29] Meta’s MTIA [18]) enable full programmer/compiler control over memory operations (when and what to read/write, when and what to evict, etc). AI accelerators typically employ scratchpad memory-based systems to support deterministic low-level activities [21, 36, 57, 37]. AI accelerators are also integrated into edge hardware (e.g., NPU in Apple Vision Pro’s M2 chip [2]).
Scheduling on MCM AI Accelerators. To scale with the rising compute demands of modern AI workloads, multi-chip modules (MCMs) have emerged as viable approach enabling the integration of composable, small functional dies (chiplets) on the package level to build a larger system, where they are connected together via on-package links typically through silicon interposer or organic substrates to create a network-on-package (NoP) [70, 30, 5]. Through enabling scalability via adjusting the number of chiplets on the package, as well as low verification costs [50], many chiplet-based systems have been developed for scalable deep learning inference: Simba [64] is one comprising 36 chiplets, each containing 16 processing engines to deliver up to 128 TOPs computing capability. Another example is Tesla’s DOJO chiplet-based architecture capable of scaling to exaFLOP supercomputers for large-scale machine learning. The scaling in chiplet sizes, architectures, and computational capabilities has enabled support for serving multi-model workloads together on the same MCM system with a finer degree of scheduling granularity (operator, tiles). However, multi-model schedulers face new challenges compared to their single-model counterparts considering the increased memory footprints, bandwidth contention, etc.
II-C Motivational Example
Figure 2: Motivational study on a $2\times 2$ MCM AI accelerator using batch size of 1 for 3 layers from the second ResNet-50 block and the first feed forward layer from GPT-2. Each chiplet has 4096 PEs and 10 MB L2 shared memory (Full from Section V). NN-baton [68] considers partitioning computation across chiplets only when not enough resources exist.
We consider the NN-baton [68] as our baseline scheduler as it targets scheduling single model workloads on multi-chiplet accelerators. NN-baton proposes to partition a single model workload across several chiplets whenever its computational demands exceeds a single chiplet’s capacity, and employs a unified dataflow across the chiplets. As heterogeneous accelerators proliferate [37, 77], chiplets technology has facilitated their integration on the package level. Consider a small heterogeneous 2$\times$2 MCM containing 3 NVDLA-like (weight stationary) and 1 Shidiannao-like (output stationary accelerators, and consider a small multi-model workload constituting 3 layers from the second ResNet-50 block and one fully connected layer from GPT-L. We analyze the schedules yielded through NN-baton and our scheduler as follows.
Single model case. We show the single model scheduling results for the ResNet-50 workload in Figure 2 [A1-A3]. As each chiplet possess sufficient resources to process the ResNet-50 workload, NN-baton schedules the workload onto a single chiplet. As shown, scheduling the ResNet-50 workload to the NVDLA-like chiplet (A2) experiences 0.78$\times$ the EDP as that from the Shidiannao-like chiplet (A1). However, a more nuanced schedule (A3) identified through our scheduler leverages heterogeneity by distributing the ResNet-50 layers across the heterogeneous chiplets, sustaining 0.52$\times$ less EDP than (A1) through catering to individual layer affinities.
Multi-model case., In Figure 2 (B1-B3), NN-baton (B1) is agnostic to the heterogeneous MCM composition, executing each model workload sequentially on its starting chiplet 1. We show two schedules that are sampled through our schedules: (i) In (B2), our scheduler can spatially distribute the ResNet-50 and GPT-2 workloads across the NVDLA-like and Shidiannao-like accelerators, respectively, leading the EDP to be 0.3$\times$ that in (B1). (ii) In (B3), Our scheduler recognizes a spatio-temporal optimization to leverage the heterogeneous chiplet pipelining for the ResNet-50 in one time window, and then schedule the GPT-2 layer on the Shidiannao layer in the following time window, leading EDP to be 0.28$\times$ that of (B1).
II-D Complexity of the multi-model scheduling space
To understand the scale of the multi-model scheduling problem, we analyze its search space complexity. Let a multi-model workload constitute $N$ models, each model containing $L_{i}$ layers, and $L$=$\sum_{i}^{N}L_{i}$. Let $C$ be the total number of accelerator chiplets on an MCM. Then, a characterization of the multi-model scheduling space can be given as $\mathcal{O}(C^{L}\times\frac{L}{L_{1}!L_{2}!\cdots L_{N}!})$.
The first term ($C^{L}$) covers the set of possible chiplet assignments for each layer (spatial complexity); whereas the multinomial coefficient ($\frac{L}{L_{1}!L_{2}!\cdots L_{N}!})$ covers the number of ways to interleave multiple sequences of layers, while maintaining the layer dependencies for each model. In the small motivational example above, this complexity accounts for a total of $\mathcal{O}(1536)$ scheduling possibilities. If we consider a more practical case involving a ResNet-50 and UNet models ($L1=50$ and $L2=23$) on a full Simba system ($C=36$), the complexity becomes $\sim$ $\mathcal{O}(10^{56})$, showcasing an exponential rise as the models grow in number and complexity.
II-E Summary on the challenges of multi-model scheduling
We summarize the unique scheduling challenges for multi-model workloads compared to their single-model counterparts:
  • Layer sequence permutations. In a single model scenario, the space of computation assignments is defined based on dependent layer sequences from a single model. Whereas in a multi-model scenario, independent layer sequences from different models also exist, compounding the decision space complexity as a result of permutations.
  • Spatial mapping conditioning. The quality of one model’s schedule is affected by the spatial mappings of other models’ mappings due to resource availability, bandwidth contention, added data travel times, and so on.
  • Heterogeneous Integration Trade-offs. Performance efficiency depends on the underlying pattern of heterogeneous integration, and the diversity within and across model workloads imply that no single pattern fits all.
@@ II-E Summary on the challenges of multi-model scheduling
To address the challenges and search complexity, one approach is to formulate the problem as a multi-level decision problem where each decision subspace is a tractable problem [37, 10]. We adopt a similar approach and formulate the MCM multi-model workload scheduling as multiple-level decision problem, as shown in Figure 3. We detail our problem formulation and performance modeling methodology next.
@@ III System Modeling and Problem Formulation
Figure 3: An overview of our scheduling algorithm for multi-model workloads on heterogeneous MCM AI accelerators.
To develop a systematic approach to navigate complex search space, we formulate the scheduling problem of multi-model workloads on a heterogeneous MCM AI accelerator.
@@ III-A Base Formulation
TABLE I: Notation used in the formulation.
Notation Description
$Sc$ Multi-model workload scenario
$m_{i}$ the i-th model from scenario $Sc$
$layer_{i,j}$ the j-th layer from model i
$H$ MCM hardware
$C$ Set of accelerator chiplets on $H$
$c_{i}$ the i-th chiplet from $C$
$DF$ set of supported dataflows on $H$
$n_{df_{i}}$ Number of chiplets adopting the i-th dataflow
$BW_{offchip}$ offchip bandwidth
$BW_{nop}$ Network-on-package bandwidth
$df$ Dataflow
$N_{PE}$ Number of processing engines
$BW_{noc}$ Network-on-chip bandwidth
$Sz_{mem}$ The memory size in $c$
$TW(Sc)$ The set of time windows for $Sc$
$tw(Sc)$ an execution time window for $Sc$ on $H$
$T_{s}$ time window start
$T_{tw}$ time window end
$L(tw(Sc))$ set of layers executable on $H$ during $tw(Sc)$
$SG$ Set of all valid segments for $tw(Sc)$
$sg(tw(Sc))$ a layer segment from $L(tw(Sc))$
$Sp_{tw}(Sc)$ Time window partioning space for $Sc$
$Sp_{sg}(tw)$ Layer segmentation space at $tw$
$SS_{TW}(tw(Sc),H)$ The scheduling space for $tw(Sc)$ on $H$
$SS_{Sc}(H)$ The overall scheduling space for $Sc$ on $H$
$sched(Sc,H)$ A scheduling instance for $Sc$ on $H$
$Lat^{i}_{j}(\mathbb{A})$ Latency evaluation for $\mathbb{A}$ given identifiers $i,j$
$E^{i}_{j}(\mathbb{A})$ Energy evaluation for $\mathbb{A}$ given identifiers $i,j$
$Sz_{data}$ Size of transmission data
$n_{hops}$ Number of hops from src to destination
$n_{splits}$ Number of time window splits
@@
Unless otherwise stated, we assume the 2D mesh topology for NoP like Simba [64], and chiplets on two sides (left and right) of the packages have off-chip interfaces.
@@ III-B Workload Partitioning Space
To develop a systematic optimization algorithm for layer segmentation i each time window, we need to define the conditions of valid layer segments, provided as follows:
@@
$\bigcup_{sg\in SG}sg=L(tw(Sc))\land[\forall sg_{i}\neq sg_{j}\in SG,sg_{i}\cap sg_{j}=\emptyset]$
@@
$\bigcup_{tw\in TW(Sc)}tw\text{=}L(Sc)\land[\forall tw_{i}\neq tw_{j}\in TW(Sc),tw_{i}\cap tw_{j}=\emptyset]$
Both Theorem 1 and  Theorem 2 indicate that the segments and time windows need to be partitions of the workload and time window layers, respectively. Combining all definitions in this section, we formulate the workload partitioning space into the time window and segment as follows:
@@ III-C Scheduling Space
=$\text{=}\{(sg,c,j)|sg\in SG\land c\in C_{H}\land j\in\mathbb{N}\land val(sg,tw(Sc))\}$
@@ III-D Scheduling Problem
The optimization metric can be chosen by users depending on the use case. In our scheduler, we adopt a comprehensive and customizable score that thoroughly considers all of latency, energy, and energy-delay product (EDP), allowing users to configure their own optimization metrics, which can be the mentioned frequently used metrics, or a user-defined function that takes a schedule instance and generates a custom metric.
III-E Performance Modeling
In order to evaluate schedules on target MCM AI accelerator hardware, we extend MAESTRO [35] to the chiplet domain and model the latency experienced on MCM AI accelerators when concurrently executing multi-model workloads We discuss our latency evaluation methodology in detail next.
@@ III-E Performance Modeling
$Lat^{comp}(l)$ is the layer computation cost dependent on chiplet parameters in Definition 2; $Lat^{ip\_com}(l)$ is latency incurred from loading the layer operands (input activations and weights); $Lat^{op\_com}(l)$ is transmission latency of output activation to a subsequent layer. $Lat^{com}$ is defined as:
@@
where communication costs are incurred when transmitting data to/from another chiplet on package or the offchip memory. The first term $\frac{Sz_{data}}{BW}$ reflects transmission latency; the second term is captures propagation latency across $n_{hops}$ between the source and destination; $\delta$ is an additional latency term for potential NoP traffic conflicts; $Lat_{mem}$ is the offchip memory read/write latency based on memory bandwidth.
@@
The first term represents the sum of individual layer computational latencies; $Lat^{ip\_com}$ is the initial external off-chiplet data transfer to load inputs and weights; $Lat^{op\_com}$ is the transmission latency from transmitting segment output data to the next segment or offchip memory. From here, we can define the time window latency as:
$Lat(tw)=\max\limits_{SG_{m}\subset SG}\;\;[Lat(SG_{m})]$
where for a time window’s assigned set of segments, $SG$, the overall window latency is the maximum latency incurred by a subset of segments, $SG_{m}\in SG$, associated with a model $m$ and necessitating sequential execution. $Lat(SG_{m})$ can be estimated as the pipelining latency across segments – that is, given model $m$ requires processing data with batch size, $b$, and the max number of samples any chiplet can process at a time is given by mini-batch size, $b^{\prime}<=b$, we get:
$Lat(SG_{m})\text{=}\sum_{sg_{k}\in SG_{m}}Lat(sg_{k}|b^{\prime})+(\frac{b}{b^{\prime}}-1)\times\max\limits_{sg_{k}}[Lat(sg_{k}|b^{\prime})]$
@@
Energy Modeling. Albeit similar to latency, energy costs are always aggregated considering communication energy costs of data sizes, $n_{hops}$, bit transmission energy, and memory access.
IV Proposed Scheduling Framework: SCAR
We discuss our scheduling framework, SCAR, for multi-model workloads on heterogeneous MCMs based on the hierarchical search space characterization and problem formulation in Section III. As illustrated in Figure 3, our scheduling algorithm is a two-level approach: top-level and per-window searches. Top-level search is responsible for selecting layers in each model to be scheduled within a time window and determining the initial number of chiplet nodes for each model. Per-window search explores the spatial and temporal partitioning (tiles or layer segments) of the layers in each model at the chiplet granularity. To explore the chiplet granularity tiling space, we generate valid inter-chiplet-pipelined schedules utilizing a scheduling tree structure inspired by the RA Tree [8]. Each schedule is evaluated using our custom heterogeneous MCM cost model which provides feedback to the chiplet level tiling (”layer segmentation” in  Figure 3) with expected metrics (latency, energy, EDP, etc.).
We codify our scheduling algorithm into a software framework, as illustrated in Figure 4. As inputs, our scheduling framework receives (1) description files of the multi-model workloads (layer parameters, topology, dependencies, etc.) and (2) a description file of the MCM hardware specification (the number of chiplets, the shape, and chiplet arrays dataflow organization, NoP bandwidth, on-chiplet memory size, etc.). As outputs, our scheduling framework reports an optimized schedule with expected metrics such as latency, energy, EDP, or other user-defined metrics as a combination of latency and energy. Our scheduling framework consists of four software engines illustrated in Figure 3 discussed next.